Grade 10Algebra

in a survey of 60 people it was found that 25 people read newspaper H, 26 read newspaper T, 26 read newspaper I, 9 read both H and I, 11 read both H and T, 8 read both T and I and 8 do not read any of the three newspapers. Find (I)number of people who read at least one of the newspaper, (ii) number of people who read all the newspaper and (iii) number of people who read only newspaper H

James John Wilson
7 Years agoGrade 10

1 Answer

Arun
7 Years ago
  1) n(H U T U I) = n(H) + n(T) + n(I) - n(H ∩ T) - n(T ∩ I) - n(I ∩ H) + n(H ∩ T ∩ I) 

==> n(H U T U I) = 25 + 26 + 26 - 11 - 8 - 9 + 3 = 52 

==> Total number of people read one or more papers is 52. 

But given survey is done among 60 people; so 8 of them do not read any of the papers mentioned. 

2) Only n(H U T) = n(H ∩ T) - n(H ∩ T ∩ I) 

==> Only n(H U T) = 11 - 3 = 8 

Similalry Only n(T U I) = 8 - 3 = 5 

and Only n(I U H) = 9 - 3 = 6 

3) Thus from the above, number of persons reading either two or three papers = 

= Only n(H U T) + Only n(T U I) + Only n(I U H) + n(H ∩ T ∩ I) = 8 + 5 + 6 + 3 = 22 

4) So number of people reading only one paper = Total number of people reading one or more papers - number reading two or three papers 

= 52 - 22 = 30 

Thus number of people reading only one paper = 30 

Alternatively, you may try to solve in another method also: 

i) Number of people read only H = Total H - (H ∩ T) - (H ∩ I) + (H ∩ T ∩ I) 
= 25 - 11 - 9 + 3 = 8 

ii) NUmber of people read only T = 26 - 11 - 8 + 3 = 10 

iii) Number of people read only I = 26 - 8 - 9 + 3 = 12 

So total reading only one paper = 8 + 10 + 12 = 30 

However of the above two, the best one to solve is with VENN Diagram, which you may try yourself.